### v2

Abstract
Weakly supervised segmentation (WSS) methods can delineate thyroid nodules in ultrasound images using training data with clinical point annotations. Existing WSS methods directly utilize the topological geometric transformations of sparse annotations as pseudo-labels, then directly compare predictions with such pseudo-labels to optimize the segmentation model. However, the noisy information contained in the single-level pseudo-labels brings low-confident references as training targets, and the incomplete metrics introduced by the single-level comparison induce low-rational learning of discriminative information between nodules and background. To address these challenges, we propose a WSS framework that generates high-confidence labels by distilling multi-level reliable prior information from clinical point annotations, and constructs a high-rationality learning strategy of comparing multi-level attributes between predictions and the proposed pseudo-labels. Specifically, our approach integrates topological geometric transformations of clinical point annotations with outputs from the MedSAM model prompted by these annotations, generating high-confidence location, foreground, and background labels. Additionally, we introduce a high-rationality learning strategy consisting of: (1) an alignment loss that measures spatial consistency between predictions and location labels, guiding the network to learn nodule feature spatial arrangements; (2) a contrastive loss that encourages feature consistency within the same region while distinguishing between foreground and background features; and (3) a prototype correlation loss that ensures consistency between feature correlations with foreground and background prototypes, refining nodule boundary delineation. Experimental results show that our method achieves state-of-the-art performance on the TN3K and DDTI datasets, proving its potential of automatically segmenting thyroid nodules in clinical practice. The code is publicly available at HCHR.
keywords:
[…]
1 Introduction
Thyroid nodule segmentation in the ultrasound image is critical for accurate thyroid disease diagnosis Tessler et al. (2017); Chen et al. (2020), but suffers from the blurred structures of anatomy with speckle noise, making it highly dependent on the expertise of the radiologist Khor et al. (2022). Employing deep learning algorithms Chen et al. (2022); Chi et al. (2023); Ozcan et al. (2024); Xiang et al. (2025) for thyroid nodule segmentation can significantly enhance diagnostic efficiency for healthcare professionals. While fully supervised algorithms Ronneberger et al. (2015); Cai et al. (2020); Zheng et al. (2021); Cao et al. (2023); Tao et al. (2022) achieve promising performance on specific datasets where precise ground truth masks are available for training, acquiring a large number of delicate annotations remains resource-intensive and time-consuming Liu et al. (2024).
Weakly supervised segmentation (WSS) algorithms offer attractive alternatives by utilizing coarse annotations such as bounding boxes Zhang et al. (2020); Mahani et al. (2022); Tian et al. (2021); Chen et al. (2024c), points Zhao and Yin (2020); Li et al. (2023b); Zhao et al. (2024b), or scribbles Wang et al. (2023); Han et al. (2024); Li et al. (2024) to achieve accurate segmentation results. WSS approaches are particularly suitable for thyroid nodule segmentation in clinical practice, since they can utilize the clinical point (aspect ratio) annotations as training supervision and obtain precise segmentation results Zhao et al. (2024b). For instance, Zhao et al. Zhao et al. (2024b) generate basic geometric shapes as radical and conservative labels for asymmetric learning, Chi et al. Chi et al. (2025) calibrated semantic features into a rational spatial distribution under the indirect, coarse guidance of the bounding box mask, and Li et al. Li et al. (2023b) generate octagon labels from point annotations as an initial contour and iteratively deform the contour.
However, existing weakly supervised methods still face the following challenges: 1) pseudo-labels generated based on topological geometric priors only are of low-confidence Zhao et al. (2024b); Chi et al. (2025); Li et al. (2023b); Zhang et al. (2020); Mahani et al. (2022); Lei et al. (2023); Fan et al. (2024), introducing label noise and potentially misleading training according to these uncertain or ambiguous conditions; 2) rigid learning strategies constructed by directly comparing the segmentation with single-level or fixed-shape segmentation pseudo-labels Zhang et al. (2020); Mahani et al. (2022); Du et al. (2023); Zhai et al. (2023), which severely limiting their flexibility and adaptability in handling diverse and complex nodule variations. It is worth note that the “level” refers to “the segmentation attributes such as location and shape or entire results” instead of “the number of pseudo-labels”.
To address the aforementioned challenges, we propose a framework where noise-free, high-confidence pseudo labels are generated at the very beginning by distilling multi-level reliable prior information from clinical point annotations, and a comprehensive, high-rationality learning strategy is constructed by comparing multi-level attributes between predictions and the proposed pseudo-labels as well. Specifically, a series of reliable attributes of weak annotations is utilized to generate multi-level initial topological shapes through geometric transformations. They are then integrated with the results obtained from MedSAM model prompted by weak annotations, leading to high-confidence multi-level pseudo-labels with over 99.5% precision in representing foreground and background regions. Furthermore, we proposed a series of loss functions to learn location-level, region-level, and boundary-level discriminative information by comparing multi-level attributes between predictions and the generated labels: 1) Alignment loss that measures the spatial projection consistency between the segmentation and the box label, guiding the network to perceive the location arrangement of nodule features; 2) Contrastive loss that reduce the intra-class variance of foreground and background features while increasing the inter-class variance between foreground and background features, guiding the network to capture the regional distribution of nodule and background features; 3) Prototype correlation loss that measures the consistency between dynamic correlation maps derived by comparing deep features with foreground and background prototypes respectively, so that the uncertain regions are gradually evolved to precise nodule edge delineation.
In summary, the contributions of our work are as follows:
  1. 1.
    We claim a series of weakly supervised segmentation objectives that precise segmentation could be obtained by learning the location and shape of nodules from partial reference. Following these objectives, high-confidence pseudo-labels and a high-rationality learning strategy are essential for improving the performance of weakly supervised thyroid nodule segmentation.
  2. 2.
    We propose a high-confidence label generation strategy that fuses geometric transformations of point annotations and segmentation provided by the prompt-guided MedSAM model. The generation of multi-level location, foreground, and background labels as learning reference prevents the misleading information from low-confidence labels during network training.
  3. 3.
    We introduce a series of high-rationality losses, including alignment, contrastive, and prototype correlation loss. These losses guide the network to capture multi-level discriminative information of thyroid nodules’ locations and shapes, significantly enhancing the reliability of the training process.
  4. 4.
    Extensive experiments on the publicly available thyroid nodule ultrasound datasets, TIN3K Gong et al. (2021) and DDTI Pedraza et al. (2015), demonstrate that our proposed framework does not only achieve state-of-the-art segmentation performance but also exhibits remarkable generalizability, making it highly adaptable to mainstream segmentation backbones.
2 Related Work
2.1 Medical Image Segmentation
Medical image segmentation serves as a cornerstone in radiology and pathology, facilitating computer-assisted automatic analysis for diagnostic and therapeutic planning Obuchowicz et al. (2024). Deep learning has driven remarkable progress in this field Rayed et al. (2024); Das et al. (2024), with fully supervised methods evolving through three key phases: (1) CNN-based Architectures: Encoder-decoder frameworks such as U-Net Ronneberger et al. (2015) and its variants (e.g., UNet++ Zhou et al. (2018), nnU-Net Isensee et al. (2021), CE-Net Tao et al. (2022)) extracted multi-scale features while preserving spatial resolution, establishing the standard for medical segmentation. (2) Transformer-enhanced Models Vaswani et al. (2017): Vision Transformers (ViTs) Dosovitskiy et al. (2021) addressed the limited long-range dependency modeling of CNNs. Networks like TransUNet Chen et al. (2024b) and Swin-Unet Cao et al. (2023) integrated self-attention mechanisms to capture global context, significantly improving accuracy for complex structures Ozcan et al. (2024); Bi et al. (2023). (3) Prompt-guided Foundation Models: Emerging paradigms like the Segment Anything Model (SAM) Kirillov et al. (2023) and its medical extension MedSAM Ma et al. (2024a) leveraged prompt engineering (e.g., bounding boxes) to enable zero-shot segmentation across diverse objects, with training on million-level or even billion-level large-scale image datasets.
Despite the architecture of the segmentation network, the learning strategy has attracted great research attention. For instance, (1) contrastive Learning optimized feature spaces by attracting positive pairs and repelling negative pairs Khosla et al. (2020); Wang et al. (2021); Zhao et al. (2021); Wang et al. (2022b). Supervised extensions Khosla et al. (2020); Wang et al. (2021) utilized segmentation masks to sample category-specific pixels, demonstrating strong boundary delineation capabilities. (2) Prototype Learning clustered intra-class feature centers to guide segmentation Snell et al. (2017); Zhou et al. (2022); Zhou and Wang (2024). Early approaches used non-learnable prototypes (e.g., class-mean features Zhou et al. (2022)) for nearest-prototype prediction. Advanced methods integrated uncertainty-aware mechanisms (e.g., entropy-based prototype fusion Lu et al. (2023)) and medical-specific adaptations like tumor sub-center prototypes for breast lesion segmentation Zhou et al. (2024).
Although providing competitive performance, current approaches faced several limitations. Firstly, fully supervised networks required extensive pixel-level labels, incurring prohibitive annotation costs Ma et al. (2024b). Secondly, the heavy reliance on prompts in prompt-guided foundation models Zhao et al. (2024a); Ma et al. (2024a) may present limitations in fully automatic settings or situations with impractical human intervention, making them unsuitable for clinical applications.
2.2 Weakly Supervised Segmentation
Weakly supervised learning represents an emerging learning paradigm that requires only a small amount of coarse-grained annotation information for model training Guo and Ge (2024); Lin et al. (2024). This approach significantly reduces the annotation workload while maintaining promising segmentation accuracy Roth et al. (2021).
Typical methods focused on directly exploiting sparse annotations or inaccurate geometric shapes to generate pseudo-labels Liu et al. (2024) for pixel-to-pixel region learning. For instance, some approaches incorporated conditional probability modeling techniques, such as conditional random fields (CRF) Zhang et al. (2020); Mahani et al. (2022), and uncertainty estimates Lei et al. (2023); Fan et al. (2024) into the training process directly using weakly supervised labels to learn predictions. Other methods generated pseudo-masks based on topological geometric transformations Wang et al. (2023); Zhao et al. (2024b); Wang and Voiculescu (2023); Li et al. (2023b). Tang et al. Tang et al. (2021) utilized ellipse masks as pseudo-labels and proposed a regional level set (RLS) loss for optimizing lesion boundary delineation.
Recently, BoxInst Tian et al. (2021) employed box annotations to localize segmentation targets, combining color similarity with graph neural networks to delineate segmentation boundaries. Inspired by BoxInst, Du et al. Du et al. (2023) proposed an algorithm that learned the location and geometric prior of organs mainly relying on the region of interest (ROI) feature, which was useful for organ segmentation with fixed prior shapes but not suitable for diverse and complex shapes.
Although recent advancements in weakly supervised segmentation methods have yielded promising results, they still relied on the quality of weak labels and the learning strategy.
2.3 Thyroid Nodule Segmentation
In recent years, algorithms for thyroid nodule segmentation based on fully supervised precise labels Chi et al. (2023); Chen et al. (2024a); Wu et al. (2024); Li et al. (2023a); Gong et al. (2023); Ma et al. (2024b) have been extensively studied. Compared to other imaging modalities like CT and MRI, ultrasound images suffer from issues such as speckle noise and low contrast, increasing the difficulty of thyroid nodule segmentation Das et al. (2024). The thyroid ultrasound image segmentation methods mainly concentrated on improving feature extraction through the innovation of a network architecture. For example, Chi et al. Chi et al. (2023) utilized transformer attention mechanisms to capture intra-frame and inter-frame contextual features in thyroid regions, yielding competitive segmentation results. Chen et al. Chen et al. (2024a) developed a multi-view learning model to encode local view features and a cross-layer graph convolution module to capture correlations between high-level and low-level features. Wu et al. Wu et al. (2024) introduced dynamic conditional encoding and a feature frequency parser based on the diffusion probabilistic model, achieving excellent segmentation results for thyroid nodules.
In order to better suit the clinical scene where delicate segmentation labels are difficult to obtain, the weakly supervision of thyroid nodule segmentation has gained great progress in recent years Yu et al. (2022); Li et al. (2023b); Zhao et al. (2024b); Chi et al. (2025). For example, Li et al. Li et al. (2023b) proposed a method that generated octagons from point annotations to serve as initial contours to iteratively refine the boundaries of thyroid nodules through active contour learning. Zhao et al. Zhao et al. (2024b) employed quadrilaterals as conservative labels and irregular ellipses as radical labels while introducing dual-branch designs to improve the consistency of pseudo-labels during training, thereby enhancing prediction accuracy. Chi et al. Chi et al. (2025) proposed a weakly supervised dual-branch learning framework to optimize the backbone segmentation network by calibrating semantic features into a rational spatial distribution under the indirect and coarse guidance of the bounding box mask.
For thyroid nodule segmentation, weak supervision has been considered as a trend with various promising models have been proposed. However, due to the quality of ultrasound images and the morphological variability of nodules, applying weak supervision to thyroid nodule segmentation in ultrasound images still faced two issues: pseudo-label noise and irrational learning strategies.
Figure 1: Overview of the proposed HCHR framework. Training process: 1) High-confidence multi-level labels are generated by fusing MedSAM results with geometric transformations as feature learning references. 2) High-rationality learning strategy is constructed through post-processing of deep features to optimize feature expression. The upper branch leverages the segmentation head to generate segmentation predictions and applies alignment loss for location-level learning. The lower branch refines feature representations for the nodule shape by calculating contrastive loss for region-level learning and the prototype correlation loss for boundary-level learning. Inference process: Images are fed into the network, followed by the segmentation head to get the results with no need for any prompts.
3 Method
3.1 Overall Framework
As shown in Fig. 1, we propose a novel weakly supervised segmentation (WSS) framework named HCHR for thyroid nodule segmentation. The framework consists of a high-confidence multi-level labels generation flow and a high-rationality multi-level learning strategy branch. In the label generation flow, we integrate prompted MedSAM results with geometric transformations of point annotations to generate high-confidence labels as training references. In the learning strategy branch, we compare multi-level attributes between predictions and generated labels by combining alignment loss, contrastive loss, and prototype correlation loss, in order to jointly learn segmentation location and delicate shape information with high rationality.
3.2 Partial Reference Segmentation Objective
Image segmentation aims to locate and delineate regions precisely. Fully supervised methods can use ground truth labels with clear location $Loc$ and shape $S$ to learn feature distributions through pixel-to-pixel comparison. However, weakly supervised approaches for thyroid ultrasound images can only refer to coarse pseudo-label cues without precise shape details, making it necessary to directly assess the rationality of feature distributions across the image domain.
Let $I$ denote the image, with $R_{f}$ as the foreground region and $R_{b}$ as the background region. Sample sets within these regions are denoted by $X_{f}$ and $X_{b}$. The network feature is represented by $\boldsymbol{\mathcal{F}}$. The subscripts $f$ and $b$ represent foreground and background. To complete the segmentation task in the image, besides the basic conditions:
1) the union of $R_{f}$ and $R_{b}$ equals the entire image ( $R_{f}\cup R_{b}=I$ );
2) their intersection is empty ( $R_{f}\cap R_{b}=\emptyset$ );
3) both regions of $R_{f}$ and $R_{b}$ are fully connected,
weakly supervised algorithms to identify regions $R_{f}$ and $R_{b}$ should also satisfy:
Objective 1: $R_{f}$ should lie within a predefined location range $Loc$, while $R_{b}$ should lie outside this range.
$R_{f}\subseteq Loc\And R_{b}\cap Loc=\emptyset.$ (1)
Objective 2: The cluster of the sample within the foreground regions $\boldsymbol{C_{(X_{f})}}$ should closely match the reference foreground cluster $\boldsymbol{C_{(R_{f})}}$, while the sample within background regions cluster $\boldsymbol{C_{(X_{b})}}$ should align with the reference background cluster $\boldsymbol{C_{(R_{f})}}$, as shown below:
$\mathcal{D}(\boldsymbol{C_{(X_{f})}},\boldsymbol{C_{(R_{f})}})<\epsilon_{f},$ (2)
$\mathcal{D}(\boldsymbol{C_{(X_{b})}},\boldsymbol{C_{(R_{b})}})<\epsilon_{b},$
where $\mathcal{D}(\cdot,\cdot)$ denotes the distance between the feature clusters and $\epsilon$ are small thresholds that ensure the proximity of the segmentation prototypes to the reference prototypes.
[…]
where $\delta$ are thresholds ensuring similarity between the sampled feature distribution and reference prototypes.
To achieve these objectives, high-confidence references used in the optimal process are essential as follows:
Reference 1: A correct range must be provided to guide the segmentation location process $Loc$.
Reference 2: High-confidence foreground and background region labels are necessary to define partial but precise foreground and background references $X_{f}$ and $X_{b}$ for image distribution $R_{f}$ and $R_{b}$ shape learning.
3.3 High-confidence Multi-level labels Generation
According to Reference 1 and Reference 2 discussed in Sec. 3.2, weakly supervised segmentation requires nodule location labels for location learning and region distribution labels for shape learning. In this section, we integrate geometric transformations of point annotations and segmentation from prompted MedSAM to generate high-confidence location labels $\boldsymbol{G_{l}}\in(0,1)^{H\times W}$ for location $Loc$ learning in Eq. (1), as well as high-confidence foreground labels $\boldsymbol{X_{f}}\in(0,1)^{H\times W}$ and background labels $\boldsymbol{X_{b}}\in(0,1)^{H\times W}$ for shape $S$ learning in Eq. (2) and Eq. (3).
Specifically, as illustrated in Fig. 1, in the high-confidence labels generation phase, we derive three geometric transformations representing low-level topological information from clinical annotations:
  • Connecting the endpoints of the annotations along each axis to form quadrilateral regions.
  • Identifying and filling the minimum bounding box enclosing these four points per target to create box regions encompassing all foreground pixels.
  • Negating the bounding box regions results in obtaining background-only regions.
The regions that contain high-level semantic information are generated by prompted MedSAM:
  • Using MedSAM with prompts computed from point annotations to obtain segmentation masks that reflect anatomical distributions from input images.
Finally, we fuse the initial topological shape ex-rectangle $\boldsymbol{g_{\text{b}}}$, in-quadrilateral $\boldsymbol{g_{\text{i}}}$, and out-rectangle $\boldsymbol{g_{\text{o}}}$ and results $\boldsymbol{y_{\text{medsam}}}$ from prompted MedSAM to generate high-confidence location $\boldsymbol{G_{l}}$, foreground $\boldsymbol{X_{f}}$, and background labels $\boldsymbol{X_{b}}$.
$\boldsymbol{G_{l}}$ $=\boldsymbol{g_{\text{b}}}\vee\boldsymbol{y_{\text{medsam}}},$ (4)
$\boldsymbol{X_{f}}$ $=\boldsymbol{g_{\text{i}}}\wedge\boldsymbol{y_{\text{medsam}}},$
$\boldsymbol{X_{b}}$ $=\boldsymbol{g_{\text{o}}}\wedge(1-\boldsymbol{y_{\text{medsam}}}),$
where $\vee$ denotes the logical OR operation applied to the positions of two masks, while $\wedge$ represents the logical AND operation.
3.4 High-rationality Multi-level Learning Strategy
According to the Objective 1, 2, and 3 discussed in Sec. 3.2, we design a high-rationality learning strategy that guides the network to learn nodule location and shape to satisfy all conditions for weakly supervised segmentation tasks, which consisting of the following three losses: alignment loss, contrastive loss and prototype correlation loss.
Figure 2: Diagram of alignment loss learning.
3.4.1 Alignment Loss for Location-level Learning
To learn the location information as described in Eq. (1), the alignment loss consists of two components: 1) alignment of projected segmentation results with location labels, and 2) topological continuity loss in high-confidence foreground regions.
As showed in Fig. 2, for a predicted result $\boldsymbol{m}\in(0,1)^{H\times W}$ and location label $\boldsymbol{G_{l}}\in(0,1)^{H\times W}$ derived from the points annotations, the directional projection Tian et al. (2021) are defined as:
$\boldsymbol{p_{x}}=\max_{col}(\boldsymbol{G_{l}})$ $,\boldsymbol{p_{y}}=\max_{row}(\boldsymbol{G_{l}}),$ (5)
$\boldsymbol{\tilde{p_{x}}}=\max_{col}(\boldsymbol{m})$ $,\boldsymbol{\tilde{p_{y}}}=\max_{row}(\boldsymbol{m}),$
where $\boldsymbol{p_{x}}\in(0,1)^{W}$ and $\boldsymbol{p_{y}}\in(0,1)^{H}$ denote projections derived from the location label $\boldsymbol{G_{l}}$ onto the x-axis and y-axis, respectively, $\boldsymbol{\tilde{p_{x}}}\in(0,1)^{W}$ and $\boldsymbol{\tilde{p_{y}}}\in(0,1)^{H}$ are the projections of predicted result $\boldsymbol{m}$ onto the x and y axis.
The first component $\mathcal{L}_{proj}$ of the alignment loss between the predicted mask and the location label is then computed as follows:
$\mathcal{L}_{\text{proj}}=(1-\frac{2\times|\boldsymbol{\tilde{p_{x}}}\cap\boldsymbol{p_{x}|}}{|\boldsymbol{\tilde{p_{x}}}|+|\boldsymbol{p_{x}}|})+(1-\frac{2\times|\boldsymbol{\tilde{p_{y}}}\cap\boldsymbol{p_{y}|}}{|\boldsymbol{\tilde{p_{y}}}|+|\boldsymbol{p_{y}}|}).$ (6)
This projection loss provides region localization constraints with high feasibility and efficiency, but still faces issues such as segmentation results lying along the projection axis or containing holes and discontinuities in the topological shape. To solve these issues, we introduce topological continuity loss as the second component of the alignment loss. The topological continuity loss $\mathcal{L}_{\text{topo}}$ is defined as:
$\mathcal{L}_{\text{topo}}=-\left[\boldsymbol{m^{\prime}}\log(\boldsymbol{X_{f}})+(1-\boldsymbol{m^{\prime}})\log(1-\boldsymbol{X_{f}})\right],$ (7)
where $\boldsymbol{m^{\prime}}={\boldsymbol{m}}\wedge\boldsymbol{X_{f}}$ denotes predicted regions ${\boldsymbol{m}}$ covered by the high-confidence foreground $\boldsymbol{X_{f}}$.
The topological continuity loss ensures topological continuity within predicted foreground regions covered by the high-confidence foreground, where foreground pixels’ precision of over $99.65\%$ as calculated in Table 1. This specific loss function is exclusively utilized to identify foreground regions to prevent local optima in coordinate axis-based optimization.
By combining loss $\mathcal{L}_{\text{proj}}$ and loss $\mathcal{L}_{\text{topo}}$, we obtain the final alignment loss $\mathcal{L}_{\text{alignment}}$ as follows:
$\mathcal{L}_{\text{alignment}}=\mathcal{L}_{\text{proj}}+\mathcal{L}_{\text{topo}}.$ (8)
This loss allows weak supervision labels to constrain segmentation localization effectively while avoiding over-fitting to a single shape.
3.4.2 Contrastive Loss for Region-level Shape Learning
Figure 3: Diagram of contrastive loss learning.
As outlined in Sec. 3.2, we propose to learn the segmentation shape by refining the feature representation of the foreground and background regions to achieve Eq. (2). Due to the features extracted by the network being inherently uncertain, multi-scale contrastive loss is exploited to optimize the inter-class variance and intra-class similarity of features sampled from the foreground and background label regions, ultimately refining the predicted foreground and background feature cluster extracted from the whole image domain.
Given high-confidence foreground areas $\boldsymbol{X_{f}}$ and background areas $\boldsymbol{X_{b}}$. As shown in Fig. 3, the generic contrastive loss with a sampling scale size $k\times k$, denoted as
$\mathcal{L}_{\text{cnt}}^{k}=\frac{1}{n}\sum_{\boldsymbol{q_{k}}^{+}\in\boldsymbol{Q_{p}}}-\log\left(\frac{\exp\left(\boldsymbol{q_{k}}\cdot\boldsymbol{q_{k}^{+}}/\tau\right)}{\exp\left(\boldsymbol{q_{k}}\cdot\boldsymbol{q_{k}^{+}}/\tau\right)+\sum_{\boldsymbol{q_{k}^{-}}\in\boldsymbol{Q_{n}}}\exp\left(\boldsymbol{q_{k}}\cdot\boldsymbol{q_{k}^{-}}/\tau\right)}\right),$ (9)
where $\boldsymbol{Q_{p}}$ and $\boldsymbol{Q_{n}}$ represent the sets of positive and negative feature embedding queues extracted from foreground and background feature map regions, respectively. $\boldsymbol{q_{k}^{+}}$ denotes a foreground feature sample of size $k\times k$, while $\boldsymbol{q_{k}^{-}}$ represents corresponding background feature samples. The anchor feature queue $\boldsymbol{q}$ is selected from high-confidence foreground regions $\boldsymbol{X_{f}}$. The temperature parameter $\tau>0$ controls the slope of the loss function and its smoothness. Empirically, we evaluate the contrastive loss at scale sizes $k=3$ and $k=1$, with $\tau=0.07$.
The contrastive loss is designed to minimize the global distance between intra-class features (i.e., either both foregrounds or both backgrounds) in the embedding space and maximize the distance between inter-class features (i.e., foreground and background). Through this learning mechanism, the obtained feature representations effectively sharpen the classification boundaries between foreground and background regions.
3.4.3 Prototype Correlation Loss for Boundary-level Shape Learning
Figure 4: Diagram of prototype correlation loss learning.
To learn pixel-wisely precise segmentation shapes, building on the segmentation objective defined as Eq. (3), we extend the shape constraint to boundary-level by introducing a prototype correlation loss, to make the uncertain pixels on the classification boundary in the feature space closer to their category prototypes.
[…]
This loss comprises two components: 1) complementary consistency between feature correlations that refer to high-confidence foreground prototype and background prototype, and 2) consistency between network segmentation results and fused correlation segmentation results.
Using metric learning, we evaluate the correlation response of each position in the feature map $\boldsymbol{\mathcal{F}}$ with respect to the foreground prototypes and background prototypes, obtaining the foreground correlation response $\boldsymbol{r_{f}}$ and the background correlation response $\boldsymbol{r_{b}}$ as follows:
[…]
Since foreground and background are mutually exclusive in segmentation tasks, the foreground correlation response $\boldsymbol{\hat{r_{f}}}$ can be derived from the background prototype $\boldsymbol{r_{b}}$ as follows:
$\boldsymbol{\hat{r_{f}}}=1-\boldsymbol{r_{b}}.$ (12)
The correlation map represents the similarity between image features and each prototype. The first component $\mathcal{L}_{\text{corr-fe}}$ of the prototype correlation loss reflects the complementary consistency between the feature correlations referring to high-confidence foreground and background prototypes, which is defined as follows:
$\mathcal{L}_{\text{corr-fe}}=-\frac{1}{2}\left[\boldsymbol{r_{f}}\log(\boldsymbol{\hat{r_{f}}})\right.+\left.(1-\boldsymbol{r_{f}}\log(1-\boldsymbol{\hat{r_{f}}})\right]-\frac{1}{2}\left[\boldsymbol{\hat{r_{f}}}\log(\boldsymbol{r_{f}})\right.+\left.(1-\boldsymbol{\hat{r_{f}}})\log(1-\boldsymbol{r_{f}})\right].$ (13)
The fused predictions $\boldsymbol{m_{c}}$ are obtained by balancing the foreground correlation maps $\boldsymbol{r_{f}}$ based on the foreground and those $\boldsymbol{\hat{r_{f}}}$ on the background prototypes.
$\boldsymbol{m_{c}}=\frac{\boldsymbol{r_{f}}+\boldsymbol{\hat{r_{f}}}}{2}.$ (14)
The second component $\mathcal{L}_{\text{corr-seg}}$ of the correlation loss that measures the consistency between the fused correlation map $\boldsymbol{m_{c}}$ and the segmentation prediction $\boldsymbol{m}$ is defined as:
$\mathcal{L}_{\text{corr-seg}}=1-\frac{2\times|\boldsymbol{m}\cap\boldsymbol{m_{c}}|}{|\boldsymbol{m}|+|\boldsymbol{m_{c}}|}.$ (15)
The total prototype correlation loss $\mathcal{L}_{\text{correlation}}$ is then calculated as follows:
$\mathcal{L}_{\text{correlation}}=\mathcal{L}_{\text{corr-fe}}+\mathcal{L}_{\text{corr-seg}}.$ (16)
By considering the complementary consistency of foreground and background prototype correlation, the algorithm obtained segmentation edges with low uncertainty. By directly propagating the fused segmentation boundaries to the predicted results, we achieve explicit shape learning, effectively addressing the challenge of refining the boundaries of nodule regions.
3.4.4 Overall Loss Function
By combining the abovementioned losses of learning location information and shape information, the overall loss function during the training process can be expressed as:
$\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{alignment}}+\lambda\mathcal{L}_{\text{cnt}}\pm\beta\mathcal{L}_{\text{correlation}},$ (17)
where $\mathcal{L}_{\text{all}}$ indicates the overall loss, $\mathcal{L}_{\text{alignment}}$ represents alignment loss. $\mathcal{L}_{\text{cnt}}$ represents the contrastive loss and $\mathcal{L}_{\text{correlation}}$ is the prototype correlation loss, $\lambda$ and $\beta$ are weighted parameter.
4 Experiments
4.1 Experimental Materials
4.1.1 Dataset
To evaluate the effectiveness of the proposed segmentation framework, we conduct ablation and comparative experiments on two publicly available thyroid ultrasound datasets: TN3K Gong et al. (2021) and DDTI Pedraza et al. (2015). The TN3K dataset consists of 3,494 high-resolution thyroid nodule images, following a standardized clinical split protocol with 2,879 images designated for training and 614 for testing. For the DDTI dataset, which contains 637 thyroid ultrasound scans, the data was randomly split into training and testing sets at a 4:1 ratio. Additionally, 10-fold cross-validation was applied to address the challenge of limited data while ensuring statistical reliability.
Notably, in both datasets, the ground truth annotation of thyroid nodules in ultrasound images for metrics was performed by experienced radiologists, while the point annotation was carried out by less-experienced senior medical students to simulate a real-world weakly supervised learning scenario.
4.1.2 Evaluation Metrics
We conducted a quantitative comparison using four standard segmentation evaluation metrics: Mean Intersection over Union (mIoU) Ling et al. (2024), Dice Similarity Coefficient (DSC) Ling et al. (2024); Wang et al. (2022a); Zhu et al. (2024), Hausdorff Distance at the 95th percentile (HD95) Zhu et al. (2024), and Prediction Precision Ling et al. (2024); Wang et al. (2022a). These metrics assess the segmentation’s overall accuracy, precision of segmented regions, boundary alignment, and the reliability of predicted positives, respectively. The definitions of each metric are as follows:
$\text{mIoU}=\frac{|A\cap B|}{|A\cup B|},$ (18)
$\text{DSC}=\frac{2|A\cap B|}{|A|+|B|},$ (19)
$\text{Precision}=\frac{|A\cap B|}{|A|},$ (20)
where $A$ is the predicted region and $B$ is the ground truth region.
[…]
where $A$ and $B$ represent the predicted and ground truth boundaries, and $\|p-q\|$ is the distance between points $p$ and $q$. $\inf_{q\in B}\|p-q\|$ denotes the minimum distance from a point $p\in A$ to the closest point $q\in B$, while $\sup_{p\in A}\inf_{q\in B}\|p-q\|$ represents the maximum of these minimum distances over all points in $A$. Similarly, $\sup_{q\in B}\inf_{p\in A}\|p-q\|$ is the maximum of the minimum distances from points in $B$ to their closest points in $A$.
4.1.3 Parameter Setting and Implementation
In the training process, we set a learning rate of 0.0001, a batch size of 16, and trained for 100 epochs with images resized to $256\times 256$. The network was optimized using the Adam optimizer. For comparative experiments, we adhered to the parameter configurations outlined in their respective papers. All experiments were carried out using PyTorch on an Nvidia 3090 GPU equipped with 24GB of memory.
4.2 Ablation Study
We conducted ablation studies to analyze the effectiveness of individual components in our framework. Table 2 and Table 3 provide an overview of different design strategies (Model P and Models A to E) with different pseudo-labels (represented with subscripts 1-3), together with the quantitative assessment of each model. The segmentation results together with feature heatmaps of example images across two datasets are visualized and qualified in Fig. 5, Fig. 6, Fig. 7 and Fig. 8.
4.2.1 Effectiveness of High-Confidence Labels Generation Strategy
Table 1: The precision (Mean $\pm$ STD, %) of different labels designated as foreground/background reference prototype region. The generation strategy of Topological-based denotes generating pseudo-labels by topological geometric transformation, Prompted SAM denotes generating pseudo-labels by MedSAM using point annotation as prompts, and High-confidence f/b represents our proposed high-confidence foreground/background labels.
Generation Strategy Pseudo Label TN3K DDTI
Foreground Background Foreground Background
Topological-based Ex-/Out-rectangle $73.44\pm 7.07$ $99.98\pm 0.51$ $73.64\pm 5.59$ $99.98\pm 0.11$
In-Quadrilateral $98.26\pm 3.05$ $93.59\pm 8.22$ $98.65\pm 0.98$ $93.98\pm 6.51$
Prompted MedSAM MedSAM Results $97.11\pm 4.44$ $97.70\pm 3.21$ $95.97\pm 3.66$ $97.85\pm 2.38$
Ours High-confidence f/b $\textbf{99.66}\pm\textbf{1.58}$ $\textbf{99.99}\pm\textbf{0.26}$ $\textbf{99.79}\pm\textbf{0.88}$ $\textbf{99.98}\pm\textbf{0.01}$
As shown in Table 1, the comparison of different pseudo-labels as supervision references in terms of precision, which is the ratio of correct pixel categories in pseudo-label pixels, showed that the proposed high-confidence labels can provide more reliable reference information. When high-confidence foreground/background labels served as supervision references to learn the foreground and background, they achieved higher precision (over 99.66% for foreground and 99.98% for background) than those obtained using solely topological geometric transformation (no more than 98.65% for foreground) and MedSAM results (no more than 97.11% for foreground and 97.85% for background). In two datasets, our generation of high-confidence foreground and background labels achieved precision exceeding 99.66% across domain distributions.
The region reference precision of high-confidence labels is parallel to 100% because we combine geometric topological transformation with topology priors and MedSAM prediction with anatomical information, further reducing areas of uncertainty. Therefore, they can serve as foreground and background reference $\boldsymbol{X_{f}}$ and $\boldsymbol{X_{b}}$ without misleading network learning, as outlined in Sec. 3.2
Table 2: Quantitative ablation results (Mean $\pm$ STD) of conducting models with different pseudo-label generation and network learning strategies on the TN3K Dataset. The backbone we used is U-Net. T denotes using quadrilateral regions as foreground label regions and out-of-rectangle regions as background label regions. M and H denote using prompt-generated results from the MedSAM as labels and proposed high-confidence labels, respectively. The $P_{n}$ indicates the model using corresponding pseudo-label following pixel-to-pixel learning strategy.
Model Labels Losses Metrics
$\mathcal{L}_{\text{alignment}}$ $\mathcal{L}_{\text{cnt}}$ $\mathcal{L}_{\text{correlation}}$ mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
$P_{1}$ T - - - $54.44\pm 21.25$ $67.47\pm 22.30$ $78.20\pm 29.05$ $29.42\pm 30.81$
$A_{1}$ T $62.38\pm 23.41$ $73.47\pm 23.33$ $76.79\pm 25.17$ $22.77\pm 29.49$
$B_{1}$ T $66.32\pm 22.02$ $77.04\pm 20.83$ $78.73\pm 22.18$ $20.60\pm 29.65$
$C_{1}$ T $66.89\pm 22.65$ $77.28\pm 21.63$ $77.21\pm 23.52$ $20.07\pm 28.13$
$D_{1}$ T $52.76\pm 25.05$ $65.11\pm 24.58$ $59.44\pm 29.14$ $30.87\pm 32.74$
$E_{1}$ T $\textbf{67.50}\pm\textbf{23.75}$ $\textbf{77.50}\pm\textbf{22.24}$ $\textbf{80.57}\pm\textbf{23.85}$ $\textbf{18.91}\pm\textbf{27.22}$
$P_{2}$ M - - - $62.39\pm 23.83$ $73.48\pm 23.33$ $76.79\pm 25.17$ $21.92\pm 31.13$
$A_{2}$ M $60.78\pm 23.47$ $72.16\pm 22.57$ $76.10\pm 23.76$ $22.19\pm 28.27$
$B_{2}$ M $66.17\pm 22.76$ $76.72\pm 21.80$ $76.77\pm 23.24$ $20.04\pm 28.05$
$C_{2}$ M $65.54\pm 22.52$ $76.31\pm 21.50$ $74.34\pm 23.63$ $19.91\pm 27.65$
$D_{2}$ M $53.78\pm 24.74$ $66.09\pm 24.32$ $59.43\pm 26.87$ $28.27\pm 31.41$
$E_{2}$ M $\textbf{67.25}\pm\textbf{22.71}$ $\textbf{77.51}\pm\textbf{21.82}$ $\textbf{77.37}\pm\textbf{23.24}$ $\textbf{17.94}\pm\textbf{27.13}$
$P_{3}$ H - - - $52.92\pm 21.14$ $66.18\pm 22.22$ $80.27\pm 26.50$ $26.17\pm 26.97$
$A_{3}$ H $62.72\pm 23.42$ $73.91\pm 22.49$ $79.42\pm 23.30$ $21.72\pm 29.60$
$B_{3}$ H $67.49\pm 22.49$ $77.67\pm 21.58$ $78.73\pm 22.38$ $20.05\pm 29.72$
$C_{3}$ H $68.73\pm 23.68$ $78.35\pm 22.59$ $79.24\pm 23.80$ $18.09\pm 28.51$
$D_{3}$ H $61.59\pm 23.50$ $73.00\pm 22.55$ $73.32\pm 27.27$ $26.83\pm 32.13$
$E_{3}$ H $\textbf{69.30}\pm\textbf{22.38}$ $\textbf{79.10}\pm\textbf{21.37}$ $\textbf{82.49}\pm\textbf{22.69}$ $\textbf{17.20}\pm\textbf{26.97}$
Table 3: Quantitative ablation results (Mean $\pm$ STD) of conducting models with different pseudo-label generation and network learning strategies on the DDTI Dataset. The backbone we used is U-Net. T denotes using quadrilateral regions as foreground label regions and out-of-rectangle regions as background label regions. M and H denote using prompt-generated results from the MedSAM as labels and proposed high-confidence labels, respectively. The $P_{n}$ indicates the model using corresponding pseudo-label following pixel-to-pixel learning strategy.
Model Labels Losses Metrics
$\mathcal{L}_{\text{alignment}}$ $\mathcal{L}_{\text{cnt}}$ $\mathcal{L}_{\text{correlation}}$ mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
$P_{1}$ T $44.85\pm 20.77$ $58.76\pm 22.36$ $65.65\pm 31.36$ $34.49\pm 20.88$
$A_{1}$ T $53.34\pm 23.05$ $66.21\pm 22.60$ $64.50\pm 27.54$ $30.65\pm 24.81$
$B_{1}$ T $59.31\pm 21.95$ $71.68\pm 20.55$ $70.77\pm 25.68$ $26.85\pm 25.05$
$C_{1}$ T $59.70\pm 22.63$ $71.78\pm 21.34$ $70.81\pm 25.57$ $26.39\pm 24.89$
$D_{1}$ T $52.11\pm 23.25$ $65.09\pm 22.59$ $61.48\pm 28.58$ $32.57\pm 24.94$
$E_{1}$ T $\textbf{60.20}\pm\textbf{22.20}$ $\textbf{72.31}\pm\textbf{20.77}$ $\textbf{72.66}\pm\textbf{25.86}$ $\textbf{25.84}\pm\textbf{25.79}$
$P_{2}$ M $56.54\pm 23.52$ $68.87\pm 22.62$ $70.10\pm 28.14$ $27.25\pm 25.68$
$A_{2}$ M $51.52\pm 22.12$ $65.58\pm 21.64$ $63.98\pm 26.04$ $31.07\pm 27.62$
$B_{2}$ M $58.25\pm 22.25$ $70.68\pm 20.15$ $72.35\pm 24.84$ $26.49\pm 25.33$
$C_{2}$ M $58.84\pm 22.46$ $71.05\pm 20.74$ $71.60\pm 26.08$ $25.93\pm 24.63$
$D_{2}$ M $49.84\pm 23.16$ $63.04\pm 22.78$ $58.67\pm 27.33$ $33.91\pm 23.60$
$E_{2}$ M $\textbf{59.82}\pm\textbf{21.98}$ $\textbf{72.14}\pm\textbf{20.76}$ $\textbf{72.96}\pm\textbf{24.60}$ $\textbf{25.40}\pm\textbf{25.57}$
$P_{3}$ H $47.31\pm 20.41$ $61.27\pm 21.59$ $74.34\pm 29.85$ $31.92\pm 23.87$
$A_{3}$ H $53.46\pm 24.04$ $66.06\pm 23.37$ $69.12\pm 25.74$ $30.70\pm 26.22$
$B_{3}$ H $60.83\pm 23.51$ $72.43\pm 22.30$ $72.75\pm 23.38$ $24.69\pm 25.22$
$C_{3}$ H $61.42\pm 22.53$ $73.18\pm 21.22$ $72.29\pm 24.90$ $25.44\pm 26.83$
$D_{3}$ H $56.29\pm 21.70$ $69.19\pm 20.83$ $71.97\pm 26.90$ $28.13\pm 24.83$
$E_{3}$ H $\textbf{62.52}\pm\textbf{20.72}$ $\textbf{74.55}\pm\textbf{18.91}$ $\textbf{74.66}\pm\textbf{24.62}$ $\textbf{23.82}\pm\textbf{24.77}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT $P_{1}$ $P_{2}$ $P_{3}$ $E_{1}$ $E_{2}$ $E_{3}$
Figure 5: Segmentation results of using diiferent weak labels as training references. $P_{1}$, $P_{2}$ and $P_{3}$ are models using the topological transformations, prompted MedSAM results and our proposed high-confidence pseudo labels discussed in Table 1 as training references following the pixel-to-pixel fully supervised manner, while $E_{1}$, $E_{2}$ and $E_{3}$ denote models using those labels as training references following the proposed high-rationality weakly supervised manner. Row 1 and 2 are example results from the TN3K dataset, while Row 3 and 4 are example results from the DDTI dataset. Red indicates correct thyroid nodule predictions, green represents the under-segmentations of thyroid nodules, and blue shows an over-segmentations of other tissues as thyroid nodules.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT $P_{1}$ $P_{2}$ $P_{3}$ $E_{1}$ $E_{2}$ $E_{3}$
Figure 6: Feature map visualizations of segmentation results in Fig. 5. $P_{1}$, $P_{2}$, $P_{3}$, $E_{1}$, $E_{2}$ and $E_{3}$ denote the same models in Fig. 5. Row 1 and 2 are example results from the TN3K dataset, while Row 3 and 4 are example results from the DDTI dataset.
4.2.2 Effectiveness of High-confidence Labels for Network Training
The effectiveness of weak labels was firstly evaluated by using them to train the network under the pixel-to-pixel fully supervised manner. As illustrated in Fig. 5 and Fig. 6, by comparing Model $P_{1}$, $P_{2}$ and $P_{3}$, it could be observed that Model $P_{2}$ with MedSAM results as labels distinguished the nodule features clearer and segmented the nodule regions more accurate. Model $P_{3}$ with our proposed high-confidence pseudo labels didn’t show obvious advantages to the Model $P_{1}$ using the topological transformations of point annotations as labels. The quantitative results in Tabel 2 and Table 3 confirmed our qualitative observations. It could be attributed to that the total amount of information provided by the high-confidence label was less than that of the MedSAM results. However, directly learning with a fully supervised manner led to the network only focusing on the “deterministic” information represented by the label region, ignoring the equally meaningful information in the uncertain regions and resulting in under-segmentation.
In contrast, when replacing the learning strategy with our proposed weakly supervised losses, by comparing Model $E_{1}$, $E_{2}$ and $E_{3}$ in Fig. 5 and Fig. 6, it was observed that Model $E_{3}$ not only provided more accurate feature extraction and nodule segmentation than Model $E_{1}$ and $E_{2}$, but also outperformed Model $P_{2}$ by providing much less under-segmentation of nodules and over-segmentation of other tissues. Moreover, as shown in Table 2 and Table 3, Model $A_{3}$ to $E_{3}$ provided generally higher segmentation metrics than the corresponding Model $A_{1}$ to $E_{1}$ and Model $A_{2}$ and $E_{2}$. It indicates that the high-confidence information provided by the proposed high-confidence labels can enable the network to accurately capture the localization, feature clustering, and prototype differences between foreground and background during weakly supervised learning, thereby improving segmentation performance, even better than following the fully supervised manner.
Considering the performance of model with every combination of weakly supervised losses was improved by the proposed high-confidence labels (Model $A_{3}$ vs. $A_{1}$ vs. $A_{2}$, Model $B_{3}$ vs. $B_{1}$ vs. $B_{2}$, Model $C_{3}$ vs. $C_{1}$ vs. $C_{2}$, Model $D_{3}$ vs. $D_{1}$ vs. $D_{2}$, Model $E_{3}$ vs. $E_{1}$ vs. $E_{2}$), we focused on discussing the performance of Model $P_{3}$, $A_{3}$, $B_{3}$, $C_{3}$, $D_{3}$ and $E_{3}$ to evaluate the effectiveness of different losses in the following sections.
4.2.3 Effectiveness of Alignment Loss for Location Learning
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT $P_{2}$ $A_{2}$ $P_{3}$ $A_{3}$ $B_{3}$ $C_{3}$ $D_{3}$ $E_{3}$
Figure 7: Segmentation results of using diiferent loss functions as learning strategies under the high-confidence pseudo labels. $P_{2}$ denotes the model using the MedSAM results as training labels and following the pixel-to-pixel fully supervised manner. $A_{2}$ denotes the model using the same labels but using the alignment loss only as the weakly supervised learning. $P_{3}$ denotes the model using the high-confidence pseudo labels discussed in Table 1 as training references following the pixel-to-pixel fully supervised manner, while $A_{3}$, $B_{3}$, $C_{3}$, $D_{3}$ and $E_{3}$ denote models using high-confidence pseudo labels as training references following different combinations of loss functions in Table 2 and Table 3. Example images are the same as those in Fig. 5. Red indicates correct thyroid nodule predictions, green represents the under-segmentations of thyroid nodules, and blue shows an over-segmentations of other tissues as thyroid nodules.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT $P_{2}$ $A_{2}$ $P_{3}$ $A_{3}$ $B_{3}$ $C_{3}$ $D_{3}$ $E_{3}$
Figure 8: Feature map visualizations of segmentation results in Fig. 7. $P_{2}$, $A_{2}$, $P_{3}$, $A_{3}$, $B_{3}$, $C_{3}$, $D_{3}$ and $E_{3}$ denote the same models in Fig. 7. Example images are the same as those in Fig. 5.
As shown in Fig. 7 and Fig. 8, Model $A_{3}$ reduced under-segmentation of nodules and over-segmentation of tissues compared to Model $P_{3}$, and obtained regions aligned to the ground truth (GT) masks. As illustrated in Table 2 and Table 3, Model $A_{3}$ obtained a significant improvement of over 6% in mIoU and DSC across both datasets, and the decreased HD95 with the cost of only a small decrease in Precision. These results indicate that the alignment loss for location learning enhanced the ability to precisely capture target locations without misleading segmentation shapes.
When comparing Model $A_{2}$ with Model $P_{2}$ in Fig. 7 and Fig. 8, it was disappointed to observe that the segmentation from Model $A_{2}$ is worse than that from Model $P_{2}$. The 2% to 5% decrese in mIoU shown in Table 2 and Table 3 confirmed the observation. It could be attributed to that when using MedSAM to constrain position, the under-segmented labels introduce some erroneous information into the localization of nodule regions. At the same time, the results from prompted MedSAM introduce multiple shape information of nodules for training Model $P_{2}$ pixel-wisely, while the alignment loss focuses primarily on region localization and requires complementary shape losses for refinement.
In addition, by comparing Model $E_{3}$ with Model $D_{3}$, the improvement of all metrics was enormous. Fig. 7 and Fig. 8 further showed that without alignment loss, the segmentation of $D_{3}$ was inaccurate in terms of the shape and region alignment of the ground truth. These results indicate that the alignment loss is necessary to obtain the correct localization of nodules, allowing the shape learning losses to play their maximum roles.
Figure 9: Performance comparison of the model that combines alignment loss and contrastive loss based on HC-labels conducted on the TN3K and DDTI datasets with different weight parameters $\lambda$ of contrastive loss.
4.2.4 Effectiveness of Contrastive Loss for Semantic Shape Learning
The weight of the contrastive loss is controlled by the parameter $\lambda$. As shown in Fig. 9, experimental results supervised by proposed high-confidence labels illustrated that the Model $B_{3}$ achieved its best comprehensive performance when $\lambda$ was set to 0.8.
As shown in Fig. 7 and Fig. 8, Model $B_{3}$ outperformed Model $A_{3}$ by providing more accurate overall shapes and feature distributions of nodules. Quantitatively, as illustrated in Table 2 and Table 3, Model $B_{3}$ obtained improvements in DSC up to 6.37%, mIoU up to 7.37%, Precision up to 3.63% over Model $A_{3}$, with HD95 reduced by at least 1.67 pixels on TN3K and 3.80 pixels on DDTI. Furthermore, Model $E_{3}$ showed improvements in mIoU by over 0.57% on the TN3K dataset and nearly 0.50% on the DDTI dataset compared to the corresponding Model $C_{3}$, respectively. These results demonstrate that the incorporation of contrastive loss ensures the predicted segmentation regions closely align with the ground truth. By refining the predicted foreground and background feature clusters extracted from the whole image domain, the contrastive loss effectively improves segmentation precision.
4.2.5 Effectiveness of Prototype Correlation Loss for Semantic Shape Learning
Figure 10: Performance comparison of the model that combines alignment loss and prototype correlation loss based on HC-labels conducted on the TN3K and DDTI datasets with different weight parameters $\beta$ of prototype correlation loss.
The weight of the prototype correlation loss is controlled by the parameter $\beta$. Experimental results supervised by high-confidence labels illustrated that Model $C_{3}$ achieved its best comprehensive performance when $\beta$ was set to 0.8 on the TN3K dataset and 0.5 on the DDTI dataset, as shown in Fig. 10.
As shown in Fig. 7 and Fig. 8, Model $C_{3}$ and $E_{3}$ provided clearer and more delicate feature distributions and boundary delineations than $A_{3}$ and $B_{3}$, respectively. Moreover, the uncertain ranges of Model $C_{3}$ and $E_{3}$ on the nodule boundaries were shrunk remarkably. The qualitative observations were confirmed by the quantitative statistics in Table 2 and 3. When comparing Model $C_{3}$ with Model $A_{3}$, it achieved nearly 4% and 6% improvement in mIoU on the TN3K and DDTI dataset, along with a significant reduction in HD95 by over 2.28 and 4 pixels on TN3K and DDTI dataset. When comparing Model $E_{3}$ with Model $B_{3}$, the inclusion of prototype correlation loss also led to superior segmentation results.
The improvement in qualitative and quantitative results proves that adding prototype correlation loss as a shape constraint effectively helps the network to make the uncertain pixels on the classification boundary in the feature space closer to their category prototypes, resulting in superior segmentation regions and edges.
4.3 Comparisons with State-of-the-art Methods
We evaluated the segmentation performance of our proposed framework and compared it with several state-of-the-art weakly supervised methods: 1) SCRF Zhang et al. (2020), which combined coarse segmentation information with conditional random fields (CRF) to produce segmenation results, 2) UNCRF Mahani et al. (2022), which refined these pseudo-labels by incorporating uncertainty estimation to enhance segmentation accuracy, 3) BoxInst Tian et al. (2021), which introduced a framework use projection loss and pairwise color affinity loss to obtain segmentation results, 4) CoarseFine Chi et al. (2025), which proposed a dual-branch framework to calibrate semantic features into nodule segmentation. 5) WSDAC Li et al. (2023b), which proposed a novel weakly supervised deep active contour model for nodule segmentation, 6) S2ME Wang et al. (2023), which introduced an entropy-guided weakly supervised polyp segmentation framework, 7) IDMPS Zhao et al. (2024b), which used an asymmetric learning framework to learn nodule segmentation from conservative labels and radical labels.
Table 4: Quantitative comparison results (Mean $\pm$ STD) of different methods on the TN3K Gong et al. (2021) dataset.
Supervision Model Metrics
mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
Weakly SCRF Zhang et al. (2020) $56.68\pm 21.85$ $69.19\pm 23.00$ $63.47\pm 22.81$ $26.22\pm 29.68$
UNCRF Mahani et al. (2022) $56.27\pm 23.89$ $68.21\pm 25.25$ $66.29\pm 24.58$ $25.11\pm 29.15$
BoxInst Tian et al. (2021) $64.49\pm 23.53$ $75.27\pm 22.36$ $78.54\pm 23.91$ $20.70\pm 27.17$
CoarseFine Chi et al. (2025) $65.46\pm 24.68$ $76.14\pm 27.06$ $75.47\pm 24.36$ $20.07\pm 28.13$
WSDAC Li et al. (2023b) $61.46\pm 17.56$ $74.25\pm 17.59$ $76.81\pm 24.02$ $19.09\pm 24.69$
S2ME Wang et al. (2023) $62.76\pm 23.37$ $73.46\pm 23.41$ $75.42\pm 25.73$ $23.77\pm 28.82$
IDMPS Zhao et al. (2024b) $66.52\pm 26.64$ $75.69\pm 26.73$ $80.40\pm 25.72$ $21.48\pm 28.83$
HCHR(ours) $\textbf{69.30}\pm\textbf{22.38}$ $\textbf{79.10}\pm\textbf{21.37}$ $\textbf{82.49}\pm\textbf{22.69}$ $\textbf{17.20}\pm\textbf{26.97}$
Fully U-Net Ronneberger et al. (2015) $68.49\pm 24.34$ $78.07\pm 22.76$ $79.08\pm 24.36$ $21.13\pm 30.69$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT U-Net SCRF UNCRF Boxinst CoarseFine WSDAC S2ME IDMPS proposed
Figure 11: Qualitative comparison results on TN3K dataset. Red indicates correct thyroid nodule predictions, green represents the under-segmentations of thyroid nodules, and blue shows an over-segmentations of other tissues as thyroid nodules.
4.3.1 Comparative Results on TN3K dataset
As shown in Fig. 11, SCRF and UNCRF tended to produce box-like segmentation with significant over-segmentation. While BoxInst, CoarseFine, and IDMPS outperformed other existing methods in high-contrast scenarios (rows 2 and 3), they still struggled with under-segmentation of incomplete nodules located at image boundaries (row 1). WSDAC and S2ME frequently under-segmented thyroid nodules in images with multiple nodules or irregular nodule shapes, and S2ME completely lost its segmentation results for images with some complex backgrounds (row 2). In contrast, our proposed method achieved more consistent segmentation regions and delicate segmentation edges, while exhibited even less over- and under-segmentation than fully supervised networks using the same backbone. Quantitative results in Table 4 further showed that our framework using the U-Net backbone outperformed state-of-the-art weakly supervised methods. Specifically, it achieved an average mIOU of 69.30%, DSC of 79.10%, Precision of 82.49%, and HD95 of 17.20 pixels. These results were even better than the fully supervised U-Net, which achieved an average mIOU of 68.49%, DSC of 78.07%, Precision of 79.08%, and HD95 of 21.13 pixels.
4.3.2 Comparative Results on DDTI dataset
Table 5: Quantitative comparison results (Mean $\pm$ STD) of different methods on the DDTI Pedraza et al. (2015) dataset.
Supervision Model Metrics
mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
Weakly SCRF Zhang et al. (2020) $50.29\pm 20.38$ $64.22\pm 20.19$ $57.27\pm 24.04$ $36.88\pm 26.26$
UNCRF Mahani et al. (2022) $49.55\pm 20.67$ $63.42\pm 20.84$ $61.52\pm 25.60$ $33.39\pm 24.46$
BoxInst Tian et al. (2021) $52.53\pm 20.04$ $66.27\pm 20.15$ $64.52\pm 29.32$ $33.77\pm 20.04$
CoarseFine Chi et al. (2025) $60.44\pm 21.38$ $72.28\pm 19.56$ $72.79\pm 24.69$ $26.39\pm 24.89$
WSDAC Li et al. (2023b) $58.60\pm 18.82$ $71.78\pm 18.01$ $\textbf{76.95}\pm\textbf{26.13}$ $26.56\pm 23.96$
S2ME Wang et al. (2023) $61.35\pm 24.39$ $72.65\pm 22.81$ $73.98\pm 26.22$ $26.09\pm 28.64$
IDMPS Zhao et al. (2024b) $61.76\pm 23.11$ $74.26\pm 22.02$ $69.44\pm 26.89$ $25.21\pm 27.66$
HCHR(ours) $\textbf{62.52}\pm\textbf{20.72}$ $\textbf{74.55}\pm\textbf{18.91}$ $74.66\pm 24.62$ $\textbf{23.82}\pm\textbf{24.77}$
Fully U-Net Ronneberger et al. (2015) $58.97\pm 23.15$ $71.05\pm 21.88$ $70.83\pm 27.23$ $27.63\pm 26.81$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT U-Net SCRF UNCRF Boxinst CoarseFine WSDAC S2ME IDMPS proposed
Figure 12: Qualitative comparison results on DDTI dataset. Red indicates correct thyroid nodule predictions, green represents the under-segmentations of thyroid nodules, and blue shows an over-segmentations of other tissues as thyroid nodules.
As illustrated in Fig. 12, fully supervised networks showed significant over-segmentation in images where the foreground and background tissues were similar (rows 1 and 3). Weakly supervised algorithms, such as SCRF and UNCRF, exhibited severe over-segmentation of small targets (rows 3) and targets with complex shapes (rows 1). BoxInst and CoarseFine were suitable for simple background issues (row 2) but struggled with over- and under-segmentation when processing low-contrast images (rows 1 and 3). WSDAC performed the highest precision with minimal over-segmentation, but tended to exhibit more under-segmentation in all cases compared to other weakly supervised methods. S2ME covered most of the segmented areas, but the segmentation results are discontinuous, and it faced more severe over-segmentation issues in all cases. IDMPS provided superior segmentation results with reduced over-segmentation across images with complex shapes (rows 1), but they still exhibited gaps in low-contrast images with over-segmentation and under-segmentation (row 3). In contrast, our algorithm not only effectively reduced over-segmentation but also improved shape adaptation, surpassing fully supervised methods.
The quantitative comparison in Table 5 further confirms he qualitative observations. Our proposed method outperformed all weakly supervised approaches, achieving a 3.55% improvement in mIoU and a 3.81 pixels reduction in HD95 compared to fully supervised methods. This proved the superior segmentation performance of our framework in terms of region alignment and shape fitting.
5 Discussion
5.1 Comparison with the State-of-the-art Weakly Supervised Segmentation Methods
Weakly Supervised Segmentation (WSS) methods have attracted increasing attention due to their ability to utilize sparse annotations for generating segmentation results, thereby reducing the need for fully annotated masks. However, these methods often encounter limitations due to label noise stemming from low-confidence pseudo-labels and insufficient discriminative features extracted for diverse and complex nodule variations through rigid learning strategies. In order to further explore the reasons for ultrasound image feature perception behind segmentation results, feature maps generated by different comparison methods are visualized in Fig. 13
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT U-Net SCRF UNCRF Boxinst CoarseFine WSDAC S2ME IDMPS proposed
Figure 13: Feature map visualizations of segmentation results in Fig. 11 and Fig. 12. All the methods are the same as those in Fig. 11 and Fig. 12.
SCRF Zhang et al. (2020) and UNCRF Mahani et al. (2022) generated box-like feature distributions and predictions because they directly utilized box labels to learn segmentation results, which introduced significant inaccuracies in shape representation and misguides the training process.
Similarly, S2ME Wang et al. (2023) and IDMPS Zhao et al. (2024b) performed well on the DDTI dataset, where shape variations are minimal. However, their performance decreased on the TN3K dataset, which suffered from diverse and irregular nodule shapes. This decline in effectiveness was due to their reliance on fixed geometric pseudo-labels for shape learning, resulting in failure to capture the high-confidence discriminative features needed for accurately segmenting nodules with complex and variable shapes
BoxInst Tian et al. (2021) exhibited reasonable performance on the larger TN3K dataset but struggled with the smaller DDTI dataset due to its heavy dependence on color similarity for learning segmentation shapes. This approach was effective for high-contrast images but required a larger volume of training data. When the amount of training samples is limited, the uncertainty of feature distribution is difficult to remove.
CoarseFine Chi et al. (2025) exhibited relatively balanced performance on two datasets, but the use of rough references to refine shapes still led to evidence of over-segmentation and under-segmentation issues in its segmentation results.
WSDAC Li et al. (2023b) consistently produced under-segmentation results because it heavily relied on initial contours and image gradients, which have been proven to be less effective for images with blurred boundaries.
In contrast, our proposed method effectively addresses these challenges through two key innovations: (1) the generation of high-confidence labels to mitigate label noise and improve training stability, and (2) the introduction of a high-rationality learning strategy to capture location-level, region-level, and boundary-level features for segmentation location and shape learning. These advancements lead to more accurate localization of nodule feature arrangement, overall shape of nodule feature distribution, and delicate boundaries between nodule and background features, showing comparable or even surpassing results to those of fully supervised networks.
5.2 The Generalizability of the Proposed Weakly Supervised Segmentation Framework
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Images GT U-Net U-Net AU-Net AU-Net DUNet DUNet CENet CENet TUNet TUNet
Figure 14: Qualitative results of different extended applications on the TN3K dataset. U-Net, AU-Net, DUNet, CENet and TUNet denote the U-Net, Dense-UNet, Attention U-Net, CENet and TransUNet models in Table 6 tarined by the given ground truth masks. U-Net, AU-Net, DUNet, CENet and TUNet denote those models trained by the proposed weakly supervised learning framework. Row 1 and 2 are examples from the TN3K dataset, while Row 3 and 4 are examples from the DDTI dataset. Red indicates correct thyroid nodule predictions, green represents the under-segmentations of thyroid nodules, and blue shows an over-segmentations of other tissues as thyroid nodules.
Table 6: Quantitative results (Mean $\pm$ STD) of applying the proposed weakly supervised segmentation framework to different segmentation backbone models on the TN3K Datasett Gong et al. (2021). Symbol indicates training the model with our proposed weakly supervised segmentation framework, while symbol represents the original fully supervised learning.
Supervision Model(Backbone) Metrics
mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
Fully U-Net Ronneberger et al. (2015) $68.49\pm 24.34$ $78.07\pm 22.76$ $79.08\pm 24.36$ $21.13\pm 30.69$
Weakly U-Net Ronneberger et al. (2015) $\textbf{69.30}\pm\textbf{22.38}$ $\textbf{79.10}\pm\textbf{21.37}$ $\textbf{82.49}\pm\textbf{22.69}$ $\textbf{17.20}\pm\textbf{26.97}$
Fully Dense-UNet Cai et al. (2020) $66.61\pm 25.74$ $76.27\pm 25.19$ $77.95\pm 26.47$ $20.48\pm 28.94$
Weakly Dense-UNet Cai et al. (2020) $\textbf{68.77}\pm\textbf{23.68}$ $\textbf{78.32}\pm\textbf{22.77}$ $\textbf{79.04}\pm\textbf{23.23}$ $\textbf{18.19}\pm\textbf{28.84}$
Fully Attention U-Net Oktay et al. (2018) $69.21\pm 26.38$ $77.85\pm 25.78$ $\textbf{82.11}\pm\textbf{26.28}$ $\textbf{16.71}\pm\textbf{27.54}$
Weakly Attention U-Net Oktay et al. (2018) $\textbf{69.49}\pm\textbf{23.83}$ $\textbf{78.85}\pm\textbf{22.84}$ $81.09\pm 24.46$ $17.60\pm 27.86$
Fully CENet Tao et al. (2022) $\textbf{74.00}\pm\textbf{20.74}$ $\textbf{82.92}\pm\textbf{18.32}$ $84.68\pm 19.28$ $\textbf{14.31}\pm\textbf{23.93}$
Weakly CENet Tao et al. (2022) $73.32\pm 20.12$ $82.56\pm 17.97$ $\textbf{85.31}\pm\textbf{18.80}$ $14.83\pm 25.37$
Fully TransUNet Chen et al. (2024b) $67.13\pm 25.15$ $76.92\pm 23.14$ $79.36\pm 25.76$ $22.14\pm 29.56$
Weakly TransUNet Chen et al. (2024b) $\textbf{68.92}\pm\textbf{23.28}$ $\textbf{78.73}\pm\textbf{21.24}$ $\textbf{81.06}\pm\textbf{22.80}$ $\textbf{19.38}\pm\textbf{29.94}$
Table 7: Quantitative results (Mean $\pm$ STD) of applying the proposed weakly supervised segmentation framework to different segmentation backbone models on the DDTI Dataset Pedraza et al. (2015). Symbol indicates training the model with our proposed weakly supervised segmentation framework, while no symbol represents the original fully supervised learning.
Supervision Model(Backbone) Metrics
mIoU (%) $\uparrow$ DSC (%) $\uparrow$ Precision (%) $\uparrow$ HD95 (px) $\downarrow$
Fully U-Net Ronneberger et al. (2015) $58.97\pm 23.15$ $71.05\pm 21.88$ $70.83\pm 27.23$ $27.63\pm 26.81$
Weakly U-Net Ronneberger et al. (2015) $\textbf{62.52}\pm\textbf{20.72}$ $\textbf{74.55}\pm\textbf{18.91}$ $\textbf{74.66}\pm\textbf{24.62}$ $\textbf{23.82}\pm\textbf{24.77}$
Fully Dense-UNet Cai et al. (2020) $55.01\pm 23.53$ $67.51\pm 23.18$ $68.41\pm 27.62$ $28.36\pm 23.34$
Weakly Dense-UNet Cai et al. (2020) $\textbf{59.00}\pm\textbf{21.07}$ $\textbf{71.64}\pm\textbf{19.77}$ $\textbf{72.50}\pm\textbf{25.21}$ $\textbf{26.50}\pm\textbf{25.33}$
Fully Attention U-Net Oktay et al. (2018) $56.07\pm 23.80$ $68.56\pm 21.88$ $68.78\pm 26.83$ $28.87\pm 25.90$
Weakly Attention U-Net Oktay et al. (2018) $\textbf{58.42}\pm\textbf{23.65}$ $\textbf{70.41}\pm\textbf{23.31}$ $\textbf{69.03}\pm\textbf{26.14}$ $\textbf{26.13}\pm\textbf{24.86}$
Fully CENet Tao et al. (2022) $68.75\pm 22.40$ $78.84\pm 20.15$ $79.50\pm 23.53$ $19.88\pm 25.19$
Weakly CENet Tao et al. (2022) $\textbf{70.35}\pm\textbf{18.85}$ $\textbf{80.78}\pm\textbf{16.63}$ $\textbf{79.65}\pm\textbf{20.72}$ $\textbf{15.08}\pm\textbf{19.59}$
Fully TransUNet Chen et al. (2024b) $48.83\pm 24.38$ $61.65\pm 24.60$ $62.11\pm 29.67$ $36.13\pm 28.04$
Weakly TransUNet Chen et al. (2024b) $\textbf{51.39}\pm\textbf{22.81}$ $\textbf{64.59}\pm\textbf{22.18}$ $\textbf{62.62}\pm\textbf{28.93}$ $\textbf{32.66}\pm\textbf{24.61}$
Instead of network architecture, our proposed network focuses on the label information refinement and learning strategy optimization of weakly supervised segmentation, leading to high generalizability to different segmentation backbones.
As shown in Fig. 14, by applying the proposed high-confidence and high-rationality weakly supervised framework, different segmentation backbone models achieved comparable or even improved performance to those trained following the pixel-to-pixel fully supervised strategy. As illustrated in Table 6, on the TN3k dataset, only CENet with fully supervised training slightly outperformed our method, with a marginal increase of 0.68% in mIoU. Other networks adopting the HCHR framework exhibited higher performance with an increase in mIoU ranging from 0.28% (Attention U-Net) to 2.16% (Dense-UNet) compared to the corresponding fully supervised framework. This is despite the fully supervised network’s use of labor-intensive, more precise ground truth masks, whereas our approach relied on point annotations for weakly supervised learning. As shown in Table 7, on the DDTI dataset with a smaller number of samples, leveraging our weakly supervised algorithm outperformed all fully supervised methods across all metrics, under the same backbone architecture. Notably, it achieved a significant improvement of over 1.5% in mIoU and a reduction of more than 1.86 pixels in HD95. These results demonstrate the exceptional generalizability of our framework, which allows for the seamless replacement of feature extraction networks.
Therefore, for weakly supervised segmentation tasks that can provide multi-level labels for localization, partial foreground and background references, our framework provides a simple but effective solution for segmenting lesions with weakly supervised annotations.
5.3 Limitations and Future work
Our method illustrates strong segmentation performance on the thyroid nodule datasets, showcasing its potential for weakly supervised learning. However, there are still avenues for further enhancement.
Firstly, while our algorithm achieved a moderate inference speed of 0.0065 seconds per image, it integrated prompted MedSAM results to generate high-confidence labels for training, which required a preprocessing step for annotations. In future work, adopting more specialized and lightweight medical foundation models to generate labels could streamline this process and further reduce training complexity.
Secondly, the model introduced three loss functions to effectively constrain the learning of position and shape with two hyperparameters $\lambda$ and $\beta$ that need to be fine-tuned through ablation experiments. These hyperparameters offer potential for further optimization. Designing a solution to find the optimal parameters can further improve the performance of the algorithm in the future.
As outlined in Sec. 5.2, our framework’s feature extraction backbone was designed with adaptability in mind, allowing for seamless integration with different task requirements. Future research can explore how to tailor our framework for a range of segmentation tasks by developing task-specific feature extraction backbones, unlocking even broader applications for this method.
6 Conclusion
In this paper, we present a novel weakly supervised segmentation framework for thyroid nodule segmentation based on clinical point annotations. We clarify the segmentation objective that integrates location and shape learning to indicate the training process. Our method integrates geometric transformations with topology priors and the prompted MedSAM results with anatomical information to generate high-confidence labels. Furthermore, we propose a high-rationality learning strategy through multi-level losses. The alignment loss is for precise location learning, while contrastive and prototype correlation losses are for robust shape understanding. Experimental results demonstrate superior performance compared to state-of-the-art weakly supervised methods on the TN3K and DDTI dataset. The framework is highly versatile and can be seamlessly integrated into various feature extraction architectures, offering flexibility for diverse clinical application scenarios.
7 Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 61901098, and the Provincial Fundamental Research for Liaoning under Grant 2019JH1/10100005.